OpenStudy (pulsified333):

nvolving sets A and B. Suppose for this problem that Pr[A|B]=1/2, Pr[A]=1/5, and Pr[B′]=7/10. (1) What is Pr[B|A]? (2) What is Pr[B|A′]?

OpenStudy (misty1212):

hello again!

OpenStudy (pulsified333):

hi :)

OpenStudy (misty1212):

you want \[P(A\cap B)\]right, so you can compute the ratio you need for both

OpenStudy (pulsified333):



OpenStudy (misty1212):

\[\frac{P(A\cap B)}{P(B)}=P(A|B)\] you have two out of the three numbers you need to find \(P(A\cap B)\)

OpenStudy (pulsified333):

we have A and P[A|B]

OpenStudy (misty1212):

you got \(P(B)\) and \(P(A|B)\) so you can solve for \(P(A\cap B)\)

OpenStudy (pulsified333):

We have the complement of B

OpenStudy (pulsified333):

but B = 3/10 Right?

OpenStudy (misty1212):

yeah \[P(B)=\frac{3}{10}\]

OpenStudy (misty1212):

'cause you be sippin syrup that's why!

OpenStudy (pulsified333):

so its (3/10)*(1/2)

OpenStudy (misty1212):

hmm no

OpenStudy (misty1212):

oh yes!

OpenStudy (pulsified333):

which is 4/5

OpenStudy (anonymous):

now that is wrong

OpenStudy (misty1212):

??

OpenStudy (anonymous):

\[\frac{1}{2}\times \frac{3}{10}=\frac{3}{20}\]

OpenStudy (pulsified333):

my bad, im tired

OpenStudy (misty1212):

oh the multiplication is wrong, yeah

OpenStudy (pulsified333):

so Pr(B|A)= 3/4

OpenStudy (pulsified333):

@satellite73 how would you do (2)

OpenStudy (pulsified333):

well i know A' is 4/5

OpenStudy (anonymous):

\[P(B|A')=\frac{P(A'\cap B)}{P(A')}\]

OpenStudy (pulsified333):

but how do you Find P(A′∩B)

OpenStudy (anonymous):



OpenStudy (anonymous):

\[P(B)=P(A\cap B)\]

OpenStudy (pulsified333):

B= 3/10 and P(A∩B)=3/20

OpenStudy (pulsified333):

now what?

OpenStudy (pulsified333):

@satellite73 is the answer (3/10)/(4/5)= 3/8